the area of rectangle is 162 1/2 sq.m find its breadth if its length is 16 2/3 m

Found 2 solutions by Alan3354, addingup:
Answer by Alan3354(69443)

Divide the area by the length.
Breadth = area/length

Answer by addingup(3677)

the area of rectangle is 162 1/2 sq.m find its breadth if its length is 16 2/3 m
l × b = area
16.667 × b = 162.5
b = 162.5/16.667
b = 9.75 or 9 3/4